# STACK: Adversarial Attacks on LLM Safeguard Pipelines

### v2

@@ Abstract
Frontier AI developers are relying on layers of safeguards to protect against catastrophic misuse of AI systems. Anthropic guards their latest Claude 4 Opus model using one such defense pipeline, and other frontier developers including Google DeepMind and OpenAI pledge to soon deploy similar defenses. However, the security of such pipelines is unclear, with limited prior work evaluating or attacking these pipelines. We address this gap by developing and red-teaming an open-source defense pipeline. [^1] https://github.com/AlignmentResearch/defense-in-depth-demo, built on evaluation code by Howe et al. [2025]. First, we find that a novel few-shot-prompted input and output classifier outperforms state-of-the-art open-weight safeguard model ShieldGemma across three attacks and two datasets, reducing the attack success rate (ASR) to 0% on the catastrophic misuse dataset ClearHarm. Second, we introduce a STaged AttaCK (STACK) procedure that achieves 71% ASR on ClearHarm in a black-box attack against the few-shot-prompted classifier pipeline. Finally, we also evaluate STACK in a transfer setting, achieving 33% ASR, providing initial evidence that it is feasible to design attacks with no access to the target pipeline. We conclude by suggesting specific mitigations that developers could use to thwart staged attacks.
@@ 1 Introduction
Frontier large language models (LLMs) are steadily growing more capable, making them useful for a large and expanding range of tasks—including, unfortunately, some harmful tasks [Shevlane et al., 2023]. For example, o3-mini has demonstrated human-level persuasion abilities, and was evaluated by OpenAI as posing a “medium” risk of assisting with creating certain chemical, biological, radiological and nuclear (CBRN) hazards [OpenAI, 2025]. Similar results were found for Claude 4 Opus, with Anthropic instituting ASL-3 safeguards due to its threat capabilities [Anthropic, 2025b]. Moreover, the fact that LLMs continue to be susceptible to jailbreaks [Wei et al., 2023; Yi et al., 2024] means that these dangerous capabilities could be elicited by bad actors. AI developers’ frontier safety frameworks include commitments to prevent harms arising from misuse of their models (e.g. Anthropic, 2024; OpenAI, 2023; Google DeepMind, 2025), which are likely to require preventing jailbreaks as their models reach or surpass dangerous capability thresholds.
A key concept in safety-critical fields is defense in depth [McGuiness, 2001], also known as the swiss cheese model [Reason, 1990], where multiple defenses are layered to mitigate threats. When applied to LLM security, this approach involves layering multiple defensive components or safeguards (such as text classifiers and activation probes [Alain and Bengio, 2018]) in sequence so that a harmful request must bypass all safeguards to be successful. Defense in depth is cited as a key component of Anthropic’s Responsible Scaling policy [Anthropic, 2024], Google DeepMind’s AGI safety roadmap [Shah et al., 2025], and OpenAI’s AI Preparedness framework [OpenAI, 2023], and is already being used by Anthropic as part of the set of safeguards designed to mitigate misuse risk from Claude 4 Opus [Anthropic, 2025c].
@@ 2 Related Work
Defense-in-depth pipelines. Our work is most closely related to the work of [Sharma et al., 2025] on constitutional classifiers. Sharma et al. [2025] developed a proprietary defense-in-depth pipeline, and we develop an open-source pipeline inspired by theirs for further study. However, our key contribution lies in the evaluation, novel attack, and design recommendations, as opposed to the pipeline itself.
@@
PRP [Mangaokar et al., 2024] attacks models with output classifiers by finding a jailbreak for the classifier and a “propagation prefix” that causes the model to repeat the jailbreak. Concurrent work by Yang et al. [2025a] uses a similar idea to PRP to attack vision large language models (VLLMs). Similar to both works, our STACK attack relies on inducing the model to output a jailbreak for the output classifier. However, while PRP could be easily defeated by an input classifier blocking the suspicious query, our attack has been designed to work against models protected with both input and output classifiers via finding universal jailbreaks for each component.
@@ 3 Setting
We conceptualize a defense pipeline as being a wrapper around a target model. Formally, let $\Sigma^{*}$ be the set of all strings. Let $M$ be the set of all systems operating on strings: i.e., functions $m:\Sigma^{*}\rightarrow\Sigma^{*}$. We define a defense pipeline $p:M\rightarrow M$ as a function transforming one system into another.
@@
Attackers may have different levels of access to the components of the pipeline. Figure 2 outlines the threat models we consider; see Appendix C for more information.
@@ 3.1 Motivation
Although our primary focus is misuse, our results also shed light on the efficacy of defense-in-depth pipelines used to guard against other security threats. For example, pipelines may be used to guard against prompt injections in LLM agents [Deng et al., 2025] that would otherwise cause the agent to take actions that violate the developer’s security policy, e.g., leaking sensitive information [Greshake et al., 2023], or making arbitrary API calls [Pelrine et al., 2024]. Pipelines may also be used to defend LLM overseers that monitor another LLM as part of a control mechanism [Greenblatt et al., 2024].
Figure 2: Overview of defense-in-depth threat models. The two axes are separability of the pipeline, which is how much access the attacker gets to each defense component, and opacity of the components, which is whether the attacker gets full access to the weights (white-box) or just boolean outputs (black-box). The main threat model we consider is semi-separable black-box access, where the attacker runs a query on the pipeline, and can see which component blocked the query. For transfer, we assume separable white-box access to a proxy pipeline and inseparable black-box access to the target pipeline. For more details, see Appendix C.
@@ 3.2 Datasets and Attack Success Rate
Evaluation: We use the prompted StrongREJECT classifier to compute ASR in both cases. To distinguish between superficially compliant responses, and responses that provide genuinely useful harmful information, only responses considered maximally harmful by the StrongREJECT classifier count as successes for the attacker in our evaluations.
@@ 3.3 Constraints
We measure overrefusal on acceptable inputs by evaluating the refusal rate (RR) of our pipeline on benign queries from Llama3Jailbreaks: a large dataset of queries and responses [Bailey et al., 2025] (see Appendix B.2 for more details). To facilitate fair comparison between safeguard models, we adjust the classification thresholds for each pipeline to ensure that they have the same RR of 15%. This RR is impractically high for most applications, making our experimental conditions highly favorable to the defender. See Appendix J for more details on the threshold calculation.
@@ 4.1 Baseline Attacks
The attacks we use are Best-of-N [Hughes et al., 2024], ReNeLLM [Ding et al., 2024], and PAP [Zeng et al., 2024b]. In order to make the attack more effective for small models and simplify attack generation, we make minor modifications to Best-of-N and ReNeLLM; see Appendix D.1. Each of these attacks runs for several independent iterations on each datapoint (e.g., 40 iterations for PAP), and we consider the attack to be successful against a datapoint if any of the iterations succeed in jailbreaking the system.
@@ 4.2 Models
We use Qwen3-14B, a recent open-weight instruction-tuned model [Qwen Team, 2025], with thinking disabled, as the target model $m$ throughout this work. We build defense pipelines by augmenting the target model $m$ with open-weight safeguard models run as input and output classifiers (Section 3).
@@ 4.3 Results
We find ShieldGemma (“SGemma”) is the most effective of these guard models against PAP, and is competitive on ReNeLLM and Best-of-N. However, strikingly, we find that the Gemma 2 few-shot-prompted classifier (“Gemma2”) performs best, despite no task-specific fine-tuning. It is especially remarkable that this is the same base model, gemma-2-9b, that ShieldGemma was fine-tuned from.
@@ 5 STACK: Attacking Defense-in-Depth Pipelines
Confirm is a modification of BEAST [Sadasivan et al., 2024], an optimization-based adversarial attack, with some features from FLRT [Thompson and Sklar, 2024]. See Appendix D.2 for full details on the attack, including hyperparameters. To make Confirm find universal jailbreaks that generalize across datapoints, in each iteration we sample 16 datapoints from the dataset and use the average loss across the datapoints to compare candidate adversarial strings. For black-box attacks, the loss on one datapoint is 0 or 1 based on whether the classifier accepted or rejected the datapoint with the adversarial string applied, whereas in white-box attacks the loss is the score given by the classifier.
@@ 5.1 Black-box front-to-back STACK
The first threat model we consider is black-box, semi-separable component access in which the attacker only has black-box access to the pipeline but can infer which stage (input classifier, generative model, or output classifier) a rejected query was blocked at. For example, this is the threat model currently accessible via the Claude 4 Opus API. The key challenges with this threat model are: 1) we must jailbreak the components in order, front to back, since a input classifier refusal obscures whether the query would have succeeded against the model or output classifier; and 2) we must optimize the output classifier jailbreak through the generative model, since we do not have separate access to the output classifier. [^5] In some settings, we could use prefilling to directly query the output classifier, but we assume no access to this because prefilling could easily be disabled or subjected to additional scrutiny.
@@
Figure 4: Attack Success Rate (ASR) of PAP and front-to-back STACK on the ClearHarm dataset against the “Gemma2” pipeline. (A) through (D) are different seeds of STACK, based on different input and output jailbreaks. The best seed (A) effectively bypasses the safeguards of the “Defended” model (—–), getting close to the “Undefended” (generative model only, – – –) ASR. However, STACK damages the ability of PAP to jailbreak the generative model, performing worse than vanilla PAP (– – –) on “Undefended”. The $x$-axis shows how the ASR increases with the PAP iterations applied to rephrase each datapoint, with input and output jailbreaks fixed.
@@ 6 Limitations and Future Work
Limited classifier efficacy. Although some of the pipelines evaluated significantly improve robustness relative to an undefended model, we find that black-box attacks are effective in most cases even against state-of-the-art open-weight classifiers. This highlights the need for improved classifiers, such as through using larger models, training on more diverse datasets including synthetic data generation, or potentially ensembling multiple different classifiers. We expect our qualitative takeaways to generalize to quantitatively stronger pipelines, however we do suspect the weaknesses of existing classifiers obscure differences in performance between attack methods.
@@ 7 Discussion and Recommendations
Our results highlight limitations in existing open-weight safeguard models, with our simple few-shot-prompted classifier classifier outperforming these safeguard models. We validate that defense-in-depth pipelines can serve as an effective defense against most existing attacks. However, our staged attack STACK can bypass these pipelines. We conclude with recommendations to improve these pipelines to frustrate attacks such as ours.
Defense-in-depth works – but existing open-weight safeguard models are easily broken. Our simple few-shot-prompted classifier defeats baseline black-box attacks on ClearHarm, validating the principle of defense-in-depth. By contrast, baseline attacks like PAP and ReNeLLM successfully exploited all existing open-weight safeguard models evaluated. This highlights significant room to improve open-weight safeguards for LLM deployments.
@@
Recommendations to secure pipelines. The front-to-back attack can be prevented if the attacker cannot identify which component is currently blocking the jailbreak. Current deployments like Claude 4 Opus disclose which component blocked a query in the API response. Making refusal responses look similar regardless of whether a classifier flagged the conversation or the generative model itself refused would make attack harder. For example, classifier refusals can be generated by the target model prompted to generate a refusal. Timing side-channel attacks can be mitigated by running queries through all stages of the pipeline even if an earlier component flags it.
@@ Acknowledgments
This research was funded and supported by the UK AI Security Institute. In addition, we thank Eric Winsor for inspiring our modifications to BEAST, Xiaohan Fu for implementing PAP in our codebase, and Nikolaus Howe for early contributions to the codebase.
@@ D.1 Baseline attacks
Perturbation Ours Original [Hughes et al., 2024]
@@ D.2 Confirm hyperparameters
Confirm is an attack based on BEAST [Sadasivan et al., 2024] and FLRT [Thompson and Sklar, 2024]. Both BEAST and FLRT are optimization-based token-level attacks that find an adversarial string to append to an input over multiple iterations. Each iteration generates several candidates, picking the best canditate based on a given loss function. We optimize the attack on the CircuitBreakers dataset [Zou et al., 2024] of harmful queries and responses from Llama3 taken from Bailey et al. [2025]. Our implementation maintains a buffer of attack candidates between iterations like FLRT rather than a beam search like BEAST, and we allow more types of edit to the adversarial string in each iteration (appending, inserting, deleting, and swapping tokens; whereas BEAST only allows appending).
@@
Compute cost. These attacks are possible for a small-scale actor to launch. For the black-box front-to-back attack we spent 507 H100 GPU-hours total training all the seeds (costing around \$1000 at current cloud compute prices), and for the white-box transfer attack we spent 15 H100 GPU-hours (~\$30) total training all the seeds:

### v3

@@ Abstract
Frontier AI developers are relying on layers of safeguards to protect against catastrophic misuse of AI systems. Anthropic and OpenAI guard their latest Opus 4 model and GPT-5 models using such defense pipelines, and other frontier developers including Google DeepMind pledge to soon deploy similar defenses. However, the security of such pipelines is unclear, with limited prior work evaluating or attacking these pipelines. We address this gap by developing and red-teaming an open-source defense pipeline. [^1] https://github.com/AlignmentResearch/defense-in-depth-demo, built on evaluation code by Howe et al. [2025]. First, we find that a novel few-shot-prompted input and output classifier outperforms state-of-the-art open-weight safeguard model ShieldGemma across three attacks and two datasets, reducing the attack success rate (ASR) to 0% on the catastrophic misuse dataset ClearHarm. Second, we introduce a STaged AttaCK (STACK) procedure that achieves 71% ASR on ClearHarm in a black-box attack against the few-shot-prompted classifier pipeline. Finally, we also evaluate STACK in a transfer setting, achieving 33% ASR, providing initial evidence that it is feasible to design attacks with no access to the target pipeline. We conclude by suggesting specific mitigations that developers could use to thwart staged attacks.
@@ 1 Introduction
Frontier large language models (LLMs) are steadily growing more capable, making them useful for a large and expanding range of tasks—including, unfortunately, some harmful tasks [Shevlane et al., 2023]. For example, o3-mini has demonstrated human-level persuasion abilities, and was evaluated by OpenAI as posing a “medium” risk of assisting with creating certain chemical, biological, radiological and nuclear (CBRN) hazards [OpenAI, 2025b]. Similar results were found for Claude Opus 4, with Anthropic instituting ASL-3 safeguards due to its threat capabilities [Anthropic, 2025b]. Moreover, the fact that LLMs continue to be susceptible to jailbreaks [Wei et al., 2023; Yi et al., 2024] means that these dangerous capabilities could be elicited by bad actors. AI developers’ frontier safety frameworks include commitments to prevent harms arising from misuse of their models (e.g. Anthropic, 2024; OpenAI, 2023; Google DeepMind, 2025), which are likely to require preventing jailbreaks as their models reach or surpass dangerous capability thresholds.
A key concept in safety-critical fields is defense in depth [McGuiness, 2001], also known as the swiss cheese model [Reason, 1990], where multiple defenses are layered to mitigate threats. When applied to LLM security, this approach involves layering multiple defensive components or safeguards (such as text classifiers and activation probes [Alain and Bengio, 2018]) in sequence so that a harmful request must bypass all safeguards to be successful. Defense in depth is cited as a key component of Anthropic’s Responsible Scaling policy [Anthropic, 2024], Google DeepMind’s AGI safety roadmap [Shah et al., 2025], and OpenAI’s AI Preparedness framework [OpenAI, 2023], and is already being used by Anthropic and OpenAI to safeguard their Opus 4 and GPT-5 models [Anthropic, 2025c; OpenAI, 2025a].
@@ 2 Related Work
Defense-in-depth pipelines. Our work is most closely related to the work of Sharma et al. [2025] on constitutional classifiers. Sharma et al. [2025] developed a proprietary defense-in-depth pipeline, and we develop an open-source pipeline inspired by theirs for further study. However, our key contribution lies in the evaluation, novel attack, and design recommendations, as opposed to the pipeline itself.
@@
PRP [Mangaokar et al., 2024] attacks models with output classifiers by finding a jailbreak for the classifier and a “propagation prefix” that causes the model to repeat the jailbreak. Concurrent work by Yang et al. [2025a] uses a similar idea to PRP to attack vision language models. Similar to both works, our STACK attack relies on inducing the model to output a jailbreak for the output classifier. However, while PRP could be easily defeated by an input classifier blocking the suspicious query, our attack has been designed to work against models protected with both input and output classifiers via finding universal jailbreaks for each component.
@@ 3 Setting
We conceptualize a defense pipeline as a wrapper around a target model. Let $\Sigma^{*}$ be the set of all strings, and $M$ the set of all systems operating on strings: i.e., functions $m:\Sigma^{*}\rightarrow\Sigma^{*}$. We then define a defense pipeline $p:M\rightarrow M$ as a function that transforms one system into another.
@@
Attackers may have different levels of access to the components of the pipeline. For example, some production APIs provide logits for generated tokens, while others allow prefilling of the model response. Additionally, some model providers have exposed a moderation API that can be queried separately from the model itself, allowing attacks on separate components. Figure 2 outlines all the threat models we consider; our main results are against the realistic Black Semi-separable threat model. See Appendix C for more details on threat models.
@@ 3.1 Motivation
Although our primary focus is misuse, our results also shed light on the efficacy of defense-in-depth pipelines used to guard against other security threats. For example, pipelines may be used to guard against prompt injections in LLM agents [Deng et al., 2025] that might cause the agent to take actions that violate the developer’s security policy, e.g., leaking sensitive information [Greshake et al., 2023], or making arbitrary API calls [Pelrine et al., 2024]. Pipelines may also be used to defend LLM overseers that monitor another LLM as part of a control mechanism [Greenblatt et al., 2024].
Figure 2: Overview of defense-in-depth threat models. The two axes are separability of the pipeline—how much access the attacker gets to each defense component—and opacity of the components—whether the attacker gets full access to the weights (white-box) or just boolean outputs (black-box). The main threat model we consider is semi-separable black-box access, where the attacker runs a query on the pipeline, and can see which component blocked the query. For transfer, we assume separable white-box access to a proxy pipeline and inseparable black-box access to the target pipeline. For more details, see Appendix C.
@@ 3.2 Datasets and Attack Success Rate
Evaluation: We use the prompted StrongREJECT classifier (Appendix E) to compute ASR in both cases. To distinguish between superficially compliant responses, and responses that provide genuinely useful harmful information, only responses considered maximally harmful by the StrongREJECT classifier count as successes for the attacker in our evaluations.
@@ 3.3 Constraints
We measure overrefusal on acceptable inputs by evaluating the refusal rate (RR) of our pipeline on benign queries from Llama3Jailbreaks: a large dataset of queries and responses [Bailey et al., 2025] (see Appendix B.2 for more details). To facilitate fair comparison between safeguard models, we adjust the classification thresholds for each pipeline to ensure that they have the same RR of 15%. This RR is impractically high for most applications, making our experimental conditions highly favorable to the defender. See Appendix L for more details on the threshold calculation.
@@ 4.1 Baseline Attacks
The attacks we use are Best-of-N [Hughes et al., 2024], which adds random perturbations to the prompt; ReNeLLM [Ding et al., 2024], which rewrites the prompt and nests it inside a template; and PAP [Zeng et al., 2024b], which paraphrases the prompt with persuasion techniques. In order to make the attack more effective for small models and simplify attack generation, we make minor modifications to Best-of-N and ReNeLLM; see Appendix D.1 for more explanation, hyperparameters, and details of modifications. Each of these attacks runs for several independent iterations on each datapoint (e.g., 40 iterations for PAP), and we consider the attack to be successful against a datapoint if any of the iterations succeed in jailbreaking the system.
@@ 4.2 Models
We use Qwen3-14B, a recent open-weight instruction-tuned model [Qwen Team, 2025], with thinking disabled, as the target model $m$ throughout this work. We build defense pipelines by augmenting $m$ with open-weight safeguard models run as input and output classifiers (Section 3).
@@ 4.3 Results
We find ShieldGemma (“SGemma”) is the most effective of these guard models against PAP, and is competitive against other guard models on ReNeLLM and Best-of-N. However, strikingly, we find that the Gemma 2 few-shot-prompted classifier (“Gemma2”) performs best across the board, despite no task-specific fine-tuning. It is especially remarkable that this is the same base model, gemma-2-9b, from which ShieldGemma was fine-tuned.
@@ 5 STACK: Attacking Defense-in-Depth Pipelines
Confirm is a modification of BEAST [Sadasivan et al., 2024], an optimization-based adversarial attack, with some features from FLRT [Thompson and Sklar, 2024]. See Appendix D.2 for full details on the attack. To make Confirm find universal jailbreaks that generalize across datapoints, in each iteration we sample 16 datapoints from the dataset and use the average loss to compare candidate adversarial strings. For black-box attacks, the loss on one datapoint is 0 or 1 based on whether the classifier accepted or rejected the datapoint with the adversarial string applied, whereas in white-box attacks the loss is the score given by the classifier.
@@ 5.1 Black-box front-to-back STACK
The first threat model we consider is black-box, semi-separable component access in which the attacker only has black-box access to the pipeline but can infer at which stage (input classifier, generative model, or output classifier) a rejected query was blocked at. For example, this is the threat model currently exposed by the Claude Opus 4 API. The key challenges here are: 1) we must jailbreak the components in order, front to back, since an input classifier refusal obscures whether the query would have succeeded against the model or output classifier; and 2) we must optimize the output classifier jailbreak through the generative model, since we do not have separate access to the output classifier. [^5] In some settings, we could use prefilling to directly query the output classifier, but we assume no access to this because prefilling could easily be disabled or subjected to additional scrutiny.
@@
Figure 4: Attack Success Rate (ASR) of PAP and front-to-back STACK on the ClearHarm dataset against the “Gemma2” pipeline. (A) through (D) are different seeds of STACK, based on different input and output jailbreaks. The best seed (A) effectively bypasses the safeguards of the “Defended” model (—–), getting close to the “Undefended” (generative model only, – – –) ASR. However, STACK damages the ability of PAP to jailbreak the generative model (discussed more in Section 6), performing worse than vanilla PAP (– – –) on “Undefended”. The $x$-axis shows how the ASR increases with the PAP iterations applied to rephrase each datapoint, with input and output jailbreaks fixed. See Appendix J for the results on other models and datasets.
@@ 6 Limitations and Future Work
Limited classifier efficacy. Although some of the pipelines we evaluate significantly improve robustness relative to an undefended model, we find that black-box attacks are effective in most cases even against state-of-the-art open-weight classifiers. This highlights the need for improved classifiers, such as using larger models, training on more diverse datasets including synthetic data generation, or potentially ensembling multiple classifiers. We expect our qualitative takeaways to generalize to quantitatively stronger pipelines, however we do suspect the weaknesses of existing classifiers obscure differences in performance between attack methods.
@@ 7 Discussion and Recommendations
Our results highlight limitations in existing open-weight safeguard models, with our simple few-shot-prompted classifier classifier outperforming these safeguard models. We validate that defense-in-depth pipelines can serve as an effective defense against most existing attacks. However, our staged attack STACK can bypass these pipelines.
We also found vulnerabilities in Opus 4 [Anthropic, 2025b] and GPT-5 OpenAI [2025a] through component-wise attacks, validating that the general attack approach works against frontier model deployments (Appendix K).
We conclude with recommendations to improve these pipelines to frustrate attacks such as ours.
Defense-in-depth works—but existing open-weight safeguard models are easily broken. Our simple few-shot-prompted classifier defeats baseline black-box attacks on ClearHarm, validating the principle of defense-in-depth. By contrast, baseline attacks like PAP and ReNeLLM successfully exploited all existing open-weight safeguard models evaluated. This highlights significant room to improve open-weight safeguards for LLM deployments.
@@
Recommendations to secure pipelines. The front-to-back attack can be prevented if the attacker cannot identify which component is currently blocking the jailbreak. Current deployments like Claude Opus 4 disclose which component blocked a query in the API response, while this can be inferred from side-channels across successive queries for GPT-5. Making refusal responses look similar regardless of whether a classifier flagged the conversation or the generative model itself refused would make attack harder. For example, classifier refusals can be generated by the target model prompted to generate a refusal. Timing side-channel attacks can be mitigated by running queries through all stages of the pipeline even if an earlier component flags it.
@@ Acknowledgments
This research was funded and supported by the UK AI Security Institute. In addition, we thank Eric Winsor for inspiring our modifications to BEAST, Xiaohan Fu for implementing PAP in our codebase, and Nikolaus Howe for early contributions to the codebase. We also thank Mohammad Taufeeque, Ziwei Xu, Kellin Pelrine, and Mohan Kankanhalli for their contributions in attacking frontier models.
@@ D.1 Baseline attacks
Perturbation Ours Original (Hughes et al.)
@@ D.2 Confirm hyperparameters
Confirm is an attack based on BEAST [Sadasivan et al., 2024] and FLRT [Thompson and Sklar, 2024]. Both BEAST and FLRT are optimization-based token-level attacks that find an adversarial string to append to an input over multiple iterations. Each iteration generates several candidates, picking the best candidate based on a given loss function. We optimize the attack on the CircuitBreakers dataset [Zou et al., 2024] of harmful queries and responses from Llama3 taken from Bailey et al. [2025]. Our implementation maintains a buffer of attack candidates between iterations like FLRT rather than a beam search like BEAST, and we allow more types of edit to the adversarial string in each iteration (appending, inserting, deleting, and swapping tokens; whereas BEAST only allows appending).
@@
Compute cost. These attacks are possible for a small-scale actor to launch. For the black-box front-to-back attack we spent 507 H100 GPU-hours total training all the seeds (costing around \$1000 at current cloud compute prices), and for the white-box transfer attack we spent 15 H100 GPU-hours (˜\$30) total training all the seeds: